OpenStudy (anonymous):

can someone check my answers! :) Simplify the expressions below. Write the final product in standard form and show your work to receive full credit. 1. 2x4(4x2 + 3x + 1) 8^6 + 6x^5 2. (4x – 3)(2x2 – 7x + 1) 8x^3 – 28x^2 + 4x – 6x^2 + 21x – 3 = 8x^3 – 34x^2 + 25x – 3 3. (x2 + 4x – 3)(2x2 + x + 6) 2x^2 + 4x^3 + 24x + x^3 + 4x^2 +24x – 6x^2 – 3x – 18 = 2x^4 + 5x^3 – 2x^2 + 45x – 18 4. Write a simplified polynomial expression to represent the area of the rectangle below. I*b= (x + 5) (2x - 4) = 2x^2 – 4x + 10x – 20 = 2x^2 + 6x – 20 Do these look right? :)

OpenStudy (anonymous):

I don't feel like doing all of those, but I can tell you how to check them. Put a number in for x and solve both the original and simplified forms. If you get the same number for both. You did it right. ill do the first one for an example: 2*4(4*2+3x+1)=120 8^6 + 6(2)^5=262336 so yeah, that's a bit wrong. or maybe i misread something. Anyway, hope that helps.

OpenStudy (anonymous):

****2*4(4*2+3(2)+1)=120

OpenStudy (anonymous):

Oops there were supposed to be exponents in the first one

OpenStudy (anonymous):

2x^4(4x^2 + 3x + 1)

OpenStudy (anonymous):

Could you redo the first one with the exponents?

OpenStudy (anonymous):

736. still quite a bit off. But do you get the idea? just put in a number for x (I used 2, but anything other than zero and one should work) and solve it out.

OpenStudy (anonymous):

Yep! :)

OpenStudy (cggurumanjunath):

1.8^6 + 6x^5 U HAVE MISSED ADDING 2X^4 !

OpenStudy (cggurumanjunath):

2. IT IS CORRECT.

OpenStudy (cggurumanjunath):

4.IT IS CORRECT.

OpenStudy (cggurumanjunath):